# Characterizing topological pumping of charges in exactly solvable Rice-Mele chains of the non-Hermitian variety

### v1

@@ II.3 Charge pumping for OBC
The imaginary parts of the spectra, for some representative $\mu$-values, have been shown in Fig. 3(b). Comparing with Fig. 2(b), we find that the fluctuations of of the magnitude of the imaginary parts are very high where the BOD deviates from the expected quantized values. Comparing with Fig. 2(d), we find that the fluctuations of of the magnitude of the imaginary parts are very high where the BOD deviates from the expected quantized values.
III Summary and outlook
In this letter, we have addressed the observation of NH pumping using a periodically driven Rice-Mele model, generalized to have nonreciprocal hoppings. The system, modelled on a bipartite lattice, has only nearest-neighbour-hopping terms, which makes it possible to have closed-form analytical solutions for an open chain. This special feature allows us to determine the exact forms of the right and left eigenvectors for the OBC case, using which the Chern number can be computed in the gapped regions of the spectrum. For 1d Rice-Mele chains with cyclic time-dependence of various parameters, the charge-pumping question has been earlier addressed in Ref. [27]. However, there, the authors have demonstrated the NH puming for the PBC and OBC cases using Wannier states, which necessiates the introduction of some approximations in using the eigenstates for the OBC. We have overcome this issue by using the exact solutions available for a class of 1d NH models without longer-ranged hoppings [4, 5, 6, 7].
From our results, we have found that the BOD takes a quantized real value (in the gapped parts of the spectrum) where the imaginary parts of the energy eigenvalues do not fluctuate too much within the span of a cycle. However, how to put a mathematical bound on this is an open question, which we have not tried to answer in this letter. This needs to be addressed by asking the more fundamental question, with regard to NH systems, regarding how to meaningfully define time-evolution which would capture the correct physical behaviour. We would like to emphasize that, unlike what is claimed in Ref. [27], quantized charge-pumping is not guaranteed on using a biorthogonal basis, even on using exact eigenstates.

### v2

@@ III.2 Charge pumping for OBC
The imaginary parts of the spectra, for some representative $\mu$-values, have been shown in Fig. 3(b). Comparing with Fig. 2(d), we find that the fluctuations of of the magnitude of the imaginary parts are very high where the BOD deviates from the expected quantized values.
IV Summary and outlook
In this paper, we have addressed the observation of NH pumping using a periodically driven Rice-Mele model, generalized to have nonreciprocal hoppings. The system, modelled on a bipartite lattice, has only nearest-neighbour-hopping terms, which makes it possible to have closed-form analytical solutions for an open chain. This special feature allows us to determine the exact forms of the right and left eigenvectors for the OBC case, using which the Chern number can be computed in the gapped regions of the spectrum. For 1d Rice-Mele chains with cyclic time-dependence of various parameters, the charge-pumping question has been earlier addressed in Ref. [27]. However, there, the authors have demonstrated the NH puming for the PBC and OBC cases using Wannier states, which necessitates the introduction of some approximations in using the eigenstates for the OBC. We have overcome this issue by using the exact solutions available for a class of 1d NH models without longer-ranged hoppings [4, 5, 6, 7].
From our results, we have found that the BOD takes a quantized real value (in the gapped parts of the spectrum) where the imaginary parts of the energy eigenvalues do not fluctuate too much within the span of a cycle. However, how to put a mathematical bound on this is an open question, which we have not tried to answer in this paper. This needs to be addressed by asking the more fundamental question, with respect to NH systems, regarding how to meaningfully define time-evolution which would capture the correct physical behaviour. We would like to emphasize that, unlike what is claimed in Ref. [27], quantized charge-pumping is not guaranteed on using a biorthogonal basis, even on using exact eigenstates.